# Without Paired Labeled Data: An End-to-End Self-Supervised Paradigm for UAV-View Geo-Localization

### v1

@@ Abstract
UAV-View Geo-Localization (UVGL) aims to ascertain the precise location of a UAV by retrieving the most similar GPS-tagged satellite image. However, existing methods predominantly rely on supervised learning paradigms that necessitate annotated paired data for training, which incurs substantial annotation costs and impedes large-scale deployment. To overcome this limitation, we propose the Dynamic Memory-Driven and Neighborhood Information Learning (DMNIL) network, a lightweight end-to-end self-supervised framework for UAV-view geo-localization. The DMNIL framework utilizes a dual-path clustering-based contrastive learning architecture as its baseline to model intra-view structural relationships, enhancing feature consistency and discriminability. Additionally, a dynamic memory-driven hierarchical learning module is proposed to progressively mine local and global information, reinforcing multi-level feature associations to improve model robustness. To bridge the domain gap between UAV and satellite views, we design an information-consistent evolutionary learning mechanism that systematically explores latent correlations within intra-view neighborhoods and across cross-view domains, ultimately constructing a unified cross-view feature representation space. Extensive experiments on three benchmarks (University-1652, SUES-200, and DenseUAV) demonstrate that DMNIL achieves competitive performance against state-of-the-art supervised methods while maintaining computational efficiency. Notably, this superiority is attained without relying on paired training data, underscoring the framework’s practicality for real-world deployment. Codes will be released soon.
@@
UAV-View Geo-Localization, Unsupervised Learning, Contrastive Learning.
@@ I Introduction
UAV-View Geo-Localization (UVGL) aims to accurately infer the geographic location of a query image by retrieving reference images captured from different platforms[1]. In GPS-denied environments, aligning drone-captured near-ground perspective images with GPS-tagged satellite images enables precise geo-localization of scenes. This capability is essential for various applications, including autonomous navigation, disaster response, and surveillance. However, existing UVGL methods adopt supervised learning frameworks that require explicitly paired UAV-satellite image datasets for training[2, 3, 4]. While achieving empirical success, these approaches face two fundamental limitations: 1) The labor-intensive annotation of pixel-level or instance-level UAV-satellite pairs incurs prohibitive scalability costs[5], and 2) Vast repositories of unpaired and unlabeled cross-view imagery remain unexploited due to the dependency on supervision. More critically, when deployed in unseen geographical regions or dynamic environments (e.g., urban construction sites), supervised models necessitate exhaustive data re-annotation to adapt to domain shifts, severely hindering real-world applicability. Considering these limitations, we aim to explore an unsupervised approach to UAV-view geo-localization (UUVGL), a paradigm that leverages self-supervised learning to autonomously discover latent cross-view correlations from large-scale unannotated data.
Figure 1: (A) UReID tasks, in which intra-class variations are relatively small while inter-class differences are pronounced, facilitating robust feature discrimination.(B) UUVGL tasks, by contrast, exhibit significantly larger feature discrepancies for the same physical location due to differences in imaging conditions and viewpoints, as well as increased feature ambiguity across distinct locations arising from the spatial continuity in the captured environment.
Although the label-free UUVGL approach demonstrates promising potential, it presents substantially greater challenges compared to other unsupervised image retrieval tasks, such as unsupervised person re-identification (UReID)[6], as illustrated in Fig.1. The UUVGL paradigm must address two critical issues: 1) cross-view geometric distortions caused by significant altitude discrepancies between drone (low-altitude) and satellite (high-altitude) imaging perspectives, and 2) feature domain shifts induced by viewpoint variations. Furthermore, the inherent spatial continuity of geographical landscapes results in near-identical visual characteristics among adjacent regions, exacerbating the difficulty of discriminative location identification. As demonstrated in Fig.2, benchmark evaluations of pre-trained backbone networks on mainstream UVGL and ReID datasets reveal that initiating data relationship mining in UUVGL without annotated supervision poses considerably more challenging optimization landscapes.
@@
In recent years, unsupervised image retrieval (UIR) [14, 15, 16] has advanced significantly, offering valuable insights for UUVGL. Broadly, existing UIR methods fall into two main categories: unsupervised domain adaptation [17] and fully unsupervised approaches [18, 19, 20]. The former still relies on labeled data in the source domain, restricting its applicability in fully unsupervised scenarios. In this paper, we focus on purely unsupervised UVGL and therefore do not provide a detailed discussion of unsupervised domain adaptation, although our method can be readily extended to accommodate such settings.
Fully UIR methods are primarily divided into two categories: image style transfer[21, 22, 23] and clustering-based pseudo-label generation[24, 25, 26]. Style transfer methods employ GANs [27] to bridge domain gaps by transforming source-domain images into the target style. However, these methods heavily depend on the quality of the generated images, which can easily introduce noise and fail to mine the complex relationships between samples in the target domain. Consequently, they do not fully exploit the inter-class separability in the target domain. In UUVGL tasks, style transfer methods struggle to address the imaging and spatial distribution differences between drone and satellite images and they incur significant training time and computational costs. Clustering-based pseudo-label generation methods assign pseudo-labels using algorithms such as DBSCAN [28] and K-means [29]. Yet, their effectiveness is highly dependent on the quality of the pseudo-labels. In UUVGL tasks, the spatial continuity of geographic environments and pronounced cross-view differences make feature discrimination difficult, rendering pseudo-label generation susceptible to noise. Although confidence filtering [30] can alleviate some issues, the scarcity of high-confidence samples due to view differences increases the risk of overfitting and limits generalization. Furthermore, while cluster-contrast techniques are integrated to mitigate intra-cluster inconsistencies, they do not adequately address feature variations and dynamic changes across views, particularly under significant scale and perspective variations.
@@
To address the aforementioned issues, we propose an end-to-end self-supervised framework named Dynamic Memory-Driven and Neighborhood Information Learning(DMNIL), which achieves UVGL without the need for paired training data. Specifically, the DMNIL leverages the ConvNeXt-Tiny architecture [7] for extracting coarse-grained features, adopting a dual-path cluster-based contrastive learning method [31, 32] as the baseline. Furthermore, it integrates two key modules: Dynamic Hierarchical Memory Learning (DHML) and Information Consistency Evolution Learning (ICEL). The DHML leverages the relationship between short-term and long-term memory to adaptively model intra-class feature dynamics within the same view, thereby enhancing the consistency of dynamic feature representations. The ICEL constructs a neighborhood-driven dynamic constraint mechanism based on feature distribution, integrating neighborhood consistency constraints with mutual information optimization. This module guides both intra-view and cross-view features to form consistent representations in high-dimensional feature space, effectively capturing the latent structure and mapping relationships of cross-view features. Additionally, a pseudo-label enhancement(PLE) mechanism is introduced to improve the reliability of pseudo-labels, enabling efficient self-supervised feature learning and optimization. Ultimately, we obtain a robust representation, effectively mitigating the cross-view discrepancy under unsupervised conditions. The main contributions can be summarized as follows:
  • We propose an self-supervised end-to-end framework for UVGL that unifies feature representations through dynamic hierarchical clustering association and consistency information constraints. Without paired training data, this framework learns the latent consistency and shared feature mappings.
  • We propose a novel dynamic memory hierarchical learning module tailored for UVGL, which adaptively explores intra-class viewpoint dynamics and contextual information, reinforcing consistent intra-class representations.
  • we propose an information consistency evolution learning module that leverages neighborhood dynamic constraints and mutual information optimization to enhance shared information among features. This module dynamically establishes intra-class and inter-class neighbor learning, capturing potential mappings within and across viewpoints.
@@
  • Extensive experiments demonstrate that our framework, utilizing only a lightweight backbone, achieves state-of-the-art performance on three UVGL benchmark datasets, outperforming even certain supervised methods.
@@ II RELATED WORKS
In this section, we provide a overview of the cross-view geo-localization task and the commonly used methods in UIR, laying the theoretical foundation for UUVGL.
II-A UAV View Geo-localization
The Cross-view Geo-localization (CVGL) task[33, 34, 35] focuses on image retrieval, aiming to identify matching images from a reference image database based on a given query image. This task is particularly challenging as the query and reference images are taken from different viewpoints or sensors. In the field of CVGL, numerous approaches have been proposed to tackle difficulties associated with scenarios such as based UAV view and ground view geo-localization.
UAV view geo-localization[36, 37] has made significant strides, with feature representation being crucial for cross-view tasks. LPN[38] mined fine-grained features through local pattern partitioning, while IFSs[35] enhanced representation by integrating global and local features via a multi-branch strategy. However, challenges remain in addressing spatial discrepancies and ensuring feature consistency. CAMP[34] tackled these issues with contrastive attribute mining and position-aware partitioning, while DAC[4] used domain alignment and scene consistency constraints to improve feature consistency. MEAN[2] incorporates progressive embedding diversification, global-to-local associations, and cross-domain enhanced alignment, improving feature representation and alignment under significant viewpoint and scale variations. Although recent advancements in UVGL methods, their training processes largely rely on paired annotated datasets. This heavy dependence on labeled data significantly limits their applicability in large-scale scenarios with unlabeled data. Currently, there are no effective methods to address this limitation. Therefore, we propose an self-supervised UVGL framework designed to leverage unlabeled data, thereby reducing the reliance on extensive annotations.
II-B Ground View Geo-localization
Ground view geo-localization[39, 40] techniques have recently employed various methods to address challenges such as viewpoint variations and spatial misalignments.Early approaches, such as CVM-Net[41], used shared feature spaces to embed global descriptors, while GAN-based methods reduced visual discrepancies through view synthesis. However, these relied on fixed alignment, leading to overfitting and limited generalization. Recent methods introduced refinement and advanced architectures. TransGeo[42] employed attention-based zooming and sharpness-aware optimization to improve detail learn and generalization, while GeoDTR[43] disentangled geometric features to handle perspective shifts using a Transformer-based extractor. Hard negative mining strategies, such as those in Sample4Geo[39], further improved model discriminability by focusing on challenging negatives.
In recent years, to address the dependency on paired annotations between ground and satellite images, the literature has proposed an unsupervised ground view geo-localization method[30] that utilizes Correspondence-Free Projection (CFP) to transform ground panorama images into bird’s-eye-view (BEV) images, and applies CycleGAN[44] to reduce the visual discrepancies between ground and satellite images, while introducing a threshold-based dynamic pseudo-label updating mechanism. However, this explicit alignment approach has limitations in modeling the deep semantic relationships of cross-view features and heavily relies on the quality of CycleGAN, which may introduce artifacts and inconsistencies, affecting pseudo-label accuracy. While multi-stage training improves pseudo-label quality, it increases training time and computational cost, reducing the efficiency and practical applicability of the method. Additionally, this approach may struggle with generalization in cross-domain transfer or new scenarios.This issue motivates us to explore a fully end-to-end framework in CVGL methods, aimed at deeply mining the underlying structures and potential relationships of cross-view features while avoiding the computational resource consumption associated with multi-stage training.
II-C Unsupervised Image Retrieval
In recent years, UIR methods have garnered significant attention primarily due to their ability to eliminate reliance on large-scale manually annotated data. These methods achieve effective retrieval by exploring the inherent latent structures within the data to learn discriminative feature representations. In UIR, clustering of data features is typically employed, with cluster assignments or cluster centers used as pseudo-labels to drive network learning. Representative works include DeepCluster [45] and DeepCluster-v2 [46], which enhance feature quality through iterative optimization involving feature extraction and K-means clustering. However, the quality of pseudo-labels significantly impacts retrieval performance, especially in cross-view and cross-modal scenarios where reliable cross-modal label mappings are essential. To address this, SDPL[47] introduces the Collaborative Ranking Association (CRA) module, which combines shallow and deep features to construct cross-modal ranking relationships, extracts reliable pseudo-labels through cross-validation of ranking results, and incorporates an intra-modal ranking smoothing mechanism to reduce noise interference, thereby significantly enhancing the quality and stability of pseudo-labels.
In contrast to clustering-based approaches, contrastive learning methods typically enhance feature discrimination by maximizing the similarity between different augmented views of the same image (positive pairs) and minimizing the similarity between different images (negative pairs). Early methods such as InstDisc[48] and AMDIM[49] primarily rely on large sample dictionaries or multi-view feature interactions to learn discriminative representations. The MoCo[50] series introduces a momentum update mechanism that maintains a continuously updated feature pool during training, effectively alleviating the dependence on negative sample diversity in small-batch training. SimCLR[51] further combines large-scale small-batch parallel training with data augmentation strategies, significantly improving the performance of contrastive learning. Cluster Contrast[31] generates clusters and their centers through initial clustering and introduces image-cluster center and cluster-cluster contrastive constraints within contrastive learning. The memory module continuously updates cluster centers and sample assignments, persistently optimizing cluster boundaries and member allocations to effectively address noisy samples and distribution discrepancies between cameras. In this paper, we adopt the Cluster Contrast mechanism as our baseline to implement UUVGL tasks.
Figure 3: The pipeline of the DMNIL framework, which includes dual-stream cluster memory contrastive learning, dynamic hierarchical memory learning, and information consistency evolution modules. The dual-stream cluster memory contrastive learning captures within-view feature representations, while the dynamic memory module models feature variations under changes in viewpoint and scale. The information consistency evolution module strengthens both intra-view and cross-view feature consistency through a neighborhood learning strategy, optimizing the learning process via pseudo-labels.
@@ III METHOD
As illustrated in Fig.3, the DMNIL framework integrates dynamic hierarchical memory learning and information consistency evolution learning modules, along with a cross-domain pseudo-label optimization strategy. The framework employs a weight-shared dual-stream ConvNeXt-Tiny network as its backbone architecture, utilizing a cluster memory mechanism to perform contrastive learning on features extracted from different views, serving as the reference network. To mine the dynamic behavior of intra-class features under variations in viewpoint and scale, the framework establishes a collaborative relationship between short-term and long-term memory through the dynamic hierarchical memory learning module, enabling an in-depth analysis of the dynamic changes in local features and enhancing global consistency among features. Simultaneously, the information consistency evolution learning module leverages training image features stored in instance memory to design a feature distribution-based neighborhood dynamic constraint method, which combines local neighborhood consistency rules with mutual information optimization strategies. This guides both intra-view and cross-view features to form unified representations in high-dimensional space, further uncovering the latent relationships and mapping patterns across views. Additionally, the framework incorporates a dual-joint feature-driven pseudo-label consistency enhancement mechanism to iteratively improve the reliability and semantic accuracy of pseudo-labels.
Problem Formulation: In the UUVGL task, we are given a dataset of images taken from different platforms (e.g., drones and satellites), denoted by $\{x^{d},x^{s}\}$. Here, $x^{d}$ represents drone-view images and $x^{s}$ represents satellite-view images. Unlike the supervised setting, this task providesneither explicit labels nor paired matching data, meaning we do not know which drone-view images correspond to which satellite-view images, and there is no prior knowledge about location or semantic category. During model training, we apply random sampling to drone images and satellite images separately to simulate the absence of pairwise correspondences under this unsupervised setting. Our goal is to learn robust cross-view feature representations from large-scale unlabeled data under completely label-free and pair-free conditions by leveraging methods such as contrastive learning, clustering, or self-supervised learning. By utilizing these shared semantic space representations during the inference stage, the model can effectively retrieve and match images of the same location across different views, enabling accurate geo-localization without relying on any manual annotations or prior correspondence information.
III-A Cluster Memory Contrastive Baseline
Cluster-based contrastive learning strategies have achieved notable success in unsupervised retrieval tasks[52]. However, in cross-geo-view tasks, due to the viewpoint imaging gap and viewpoint discrepancy, single-path cluster-based contrastive learning strategies struggle to effectively address cross-view discrepancies, lack view-specific information utilization, suffer from insufficient feature alignment, and fail to fully leverage the complementarity of different views. To address these challenges, this paper introduce a dual-path cluster-based contrastive learning method as our baseline network, which leverages two independent feature streams (drone and satellite views) for contrastive learning to enhance cross-view consistency. Specifically, drone and satellite images are fed into a shared backbone network, and then processed through their respective cluster-based contrastive learning paths to learn view-specific information and shared view features. Finally, the view-specific memories are updated via a momentum update strategy.

### v2

@@ Abstract
UAV-View Geo-Localization (UVGL) aims to achieve accurate localization of unmanned aerial vehicles (UAVs) by retrieving the most relevant GPS-tagged satellite images. However, existing methods heavily rely on pre-paired UAV-satellite images for supervised learning. Such dependency not only incurs high annotation costs but also severely limits scalability and practical deployment in open-world UVGL scenarios. To address these limitations, we propose an end-to-end self-supervised UVGL method. Our method leverages a shallow backbone network to extract initial features, employs clustering to generate pseudo labels, and adopts a dual-path contrastive learning architecture to learn discriminative intra-view representations. Furthermore, our method incorporates two core modules, the dynamic hierarchical memory learning module and the information consistency evolution learning module. The dynamic hierarchical memory learning module combines short-term and long-term memory to enhance intra-view feature consistency and discriminability. Meanwhile, the information consistency evolution learning module leverages a neighborhood-driven dynamic constraint mechanism to systematically capture implicit cross-view semantic correlations, thereby improving cross-view feature alignment. To further stabilize and strengthen the self-supervised training process, a pseudo-label enhancement strategy is introduced, which refines the quality of pseudo supervision. Our method ultimately constructs a unified cross-view feature representation space under self-supervised settings. Extensive experiments on three public benchmark datasets demonstrate that the proposed method consistently outperforms existing self-supervised methods and even surpasses several state-of-the-art supervised methods. These results validate the effectiveness and generalizability of our method, highlighting its potential for real-world deployment in UVGL. Our code is available at https://github.com/ISChenawei/DMNIL.
@@
UAV-View Geo-Localization, Self-Supervised Learning, Contrastive Learning.
@@ I Introduction
UAV-View Geo-Localization (UVGL) aims to achieve high-precision localization of UAV in GPS-denied scenarios by retrieving GPS-tagged satellite images that correspond to UAV-captured near-ground view images [1]. This capability is critical for applications such as autonomous navigation, disaster rescue, and emergency response. In recent years, existing UVGL methods have primarily relied on supervised learning using strictly paired UAV-satellite images [2, 3, 4]. Although these supervised methods have achieved encouraging success, they also reveal several inherent limitations: (1) acquiring instance-level paired UAV-satellite images is highly costly [5], leading to increased training overhead; (2) large amounts of unpaired cross-view image resources remain underutilized; and (3) when applied to new scenarios, these methods often require recollection and re-pairing UAV-satellite images for training to adapt to domain shifts. These limitations significantly hinder the scalability and practical deployment of UVGL in real-world scenarios. Thereby, it is necessary and of great significance to explore a self-supervised UVGL method that can effectively mine latent cross-view correlations from large-scale unpaired UAV-satellite images.
Figure 1: (A) ReID tasks are characterized by relatively small intra-class variations and pronounced inter-class differences, which facilitates effective feature discrimination. (B) In contrast, UVGL tasks suffer from large intra-class variations and small inter-class differences, leading to significant cross-view gaps and intra-view ambiguities that hinder feature discrimination.
In recent years, self-supervised person re-identification (ReID) methods [6, 7] have achieved remarkable progress, offering crucial theoretical foundations for UVGL. Typically, These methods employ clustering algorithms [8, 9] on features extracted by a backbone network to generate pseudo-labels, which serve as the initial supervision signals for self-supervised learning. However, the quality of these pseudo-labels is heavily dependent on the discriminative capability of backbone network. As illustrated in Fig. 1, compared with conventional ReID, UVGL faces additional challenges. These factors make it more difficult for the backbone to extract discriminative features, as illustrated in Fig. 2. When pseudo-labels contain a high error rate and the model lacks the capacity to capture intra-view and cross-view consistency and discriminative features, the self-supervised learning process becomes prone to catastrophic negative optimization.
@@
In this paper, we propose a novel end-to-end self-supervised UVGL method. A lightweight, shallow ConvNeXt-Tiny backbone network [10] is employed to extract features, which are subsequently clustered to generate pseudo-labels. Combined with a dual-path contrastive learning strategy [17, 18], these components form the baseline network of our self-supervised method for UVGL to learn intra-view feature representations. Building upon this baseline, we proposed the dynamic hierarchical memory learning (DHML) module and the information consistency evolution learning (ICEL) module as two key components of our method. The DHML module captures the relationship between short-term and long-term memory to learn the discriminability of intra-view feature representations. The ICEL module constructs a neighborhood-driven dynamic constraint mechanism based on feature distributions. By integrating neighborhood consistency constraints with mutual information optimization, it effectively aligns cross-view feature representations within a shared embedding space, promoting both consistency and discriminative learning. Additionally, to improve the reliability of pseudo-labels, we introduce a pseudo-label enhancement (PLE) strategy, which contributes to a more stable and effective self-supervised feature learning process. The main contributions can be summarized as follows:
  • We propose an end-to-end self-supervised method for UVGL. Built upon a shallow backbone and a dual-path contrastive learning strategy, our method integrates dynamic hierarchical memory learning and information consistency evolution learning to effectively capture and enhance feature consistency and discriminability for both intra-view and cross-view representation learning.
  • We propose a dynamic memory hierarchical learning module that captures the relationships among different levels of memory to adaptively model intra-class and inter-class feature dynamics within the intra-view, thereby enhancing the discriminability of feature representations.
  • we propose a novel information consistency evolution learning module which combines neighborhood consistency and mutual information optimization to learn consistent and discriminative cross-view representations.
@@
  • Extensive experiments demonstrate that our method, utilizing only a lightweight backbone, achieves state-of-the-art performance on three UVGL benchmarks, outperforming even certain supervised methods.
The remainder of this paper is organized as follows. Section II systematically reviews prior research. In Section III, the proposed self-supervised UVGL method is presented in detail. The experimental results are reported and analyzed in Section IV. Finally, conclusions are outlined in Section V.
@@ II RELATED WORKS
This section establishes a systematic and comprehensive theoretical framework to support the development of the proposed self-supervised UVGL method. Although this work focuses on UVGL as a sub-task of cross-view geo-localization (CVGL), we first review the overall progress in supervised CVGL. We then discuss recent advances in self-supervised CVGL methods. Finally, we summarize main methodes in self-supervised ReID.
II-A Supervised Cross-View Geo-localization
CVGL aims to achieve high-precision localization of query images by retrieving GPS-tagged satellite images. Early research in CVGL primarily focused on the ground-view CVGL [19, 20, 21], leading to the development of several benchmark datasets, such as CVUSA [22], CVACT [23], and VIGOR [24]. To address the challenge of ground-view CVGL, researchers have proposed various methods. CVM-Net [25] employed NetVLAD [26] to encode images as global descriptors, thereby reducing visual discrepancies caused by viewpoint variations. TransGeo [27] utilized attention-based zooming and sharpness-aware optimization to enhance detail learning and improve model generalization. GeoDTR [20] leveraged a Transformer-based [28] feature extractor to disentangle geometric features, effectively addressing perspective shifts. Sample4Geo [29] introduced a hard negative mining strategy, focusing on challenging negative samples to further improve discriminative capability.
With the continuous development and application of drone technology, UVGL has gradually become an important research direction and achieves significant progress [30, 31, 32]. Several benchmark datasets for UVGL have been proposed, such as University-1652[12], SUES-200, [33] and DenseUAV [34], which have driven further research in this field. In this scenario, a variety of methods have been proposed by researchers. LPN [35] mined fine-grained features through local pattern partitioning, improving the ability to capture local details. IFS [36] utilized a multi-branch strategy to effectively fuse global and local features, enhancing the diversity and robustness of feature representations. CAMP [30] improved cross-view matching accuracy by contrastive attribute mining and position-aware partitioning, while DAC [4] enhanced feature consistency and matching accuracy by introducing domain alignment and scene consistency constraints. MEAN [2] adopted progressive embedding diversification, global-to-local associations, and cross-domain enhanced alignment to further improve feature representation and alignment.
Nevertheless, the remarkable performances exhibited by these UVGL methods require extensive pre-processed UAV-satellite image pairs. In this work, we pivot our focus to the realm of self-supervised UVGL, where prior pairing relationships are absent, presenting important applications for real-world UVGL deployments.
II-B Self-Supervised Cross-View Geo-localization
Due to the dependence of CVGL on extensive pre-processed image pairs, which significantly increases costs and limits applications for real-world, researchers have begun to explore self-supervised CVGL methods. The work in [37] proposed a self-supervised method that leveraged correspondence-free projection to transform ground panorama images into bird’s-eye view images. It also applied CycleGAN [38] to generate fake satellite images, coupled with an additional SAFA module [39] to reduce discrepancies. However, this explicit alignment method has inherent limitations in capturing the deep semantic relationships of cross-view features and may introduce noise. Furthermore, the non-end-to-end training framework employed significantly increases training time and computational costs, reducing the efficiency and practical applicability. Additionally, existing work [5] utilized frozen foundation models (FMs) [40, 41] to extract features, and a self-supervised adapter was introduced to mitigate visual discrepancies between different point of view. However, this method heavily relied on the ability of FMs for feature extraction. When FMs fail to effectively handle viewpoint discrepancies, the performance of the adapter is also affected. To reduce reliance on the feature extractor, CDIKTNet [42] trained with a small amount of paired UAV-satellite images to optimize the initial feature distribution, and then performed clustering to generate pseudo-labels, thereby improving the quality of pseudo-labels generation. However, this method inherently depends on the feature extractor, as it still requires limited paired UAV-satellite images for effective learning.
The above methods have made meaningful exploratory attempts. However, these methods still face challenges, such as the additional computation time and overhead introduced by non-end-to-end frameworks, as well as dependence on the capability of feature extractor. In contrast, our method is fully end-to-end, eliminating the need for additional training processes and completely removing the reliance on feature extractor initialization.
II-C Self-Supervised Person Re-Identification
ReID [43, 44] aims at matching the same person image captured by non-overlapping cameras. Due to its critical role in video surveillance, ReID has attracted extensive research attention and achieved remarkable progress. However, these achievements are facilitated by extensive human-labeled data. To address this limitation, self-supervised learning methods [6, 7] have been proposed. Among them, clustering algorithms are widely employed to generate pseudo-labels, which serve as supervision signals for model training. In addition, benefiting from the introduction of instance-level contrastive learning [45] with memory banks and momentum update strategies [46], many methods [47, 48] consider each unlabeled sample as an individual class to learn discriminative instance-level feature representations. However, these instance-level features was modified largely during training, resulting in unstable representations [49]. To address this issue, recent studies [50, 51] have begun exploring cluster-level relationships to discover more stable and robust associations among pedestrian representations. Cluster Contrast [17] addresses the problem of cluster inconsistency by performing clustering and introducing a distinct representation for each cluster. Nevertheless, in cross-modal ReID scenarios, where images are captured by heterogeneous sensors, modality discrepancies hinder the effectiveness of these methods in learning modality-invariant features. ADCA[18] introduced a dual-path contrastive framework to bridge modality gaps and enhance cross-modal feature alignment. Building on this, SDCL [52] employs a dual-path Transformer for shallow-to-deep contrastive learning, and refines pseudo-labels via consistency constraints across feature hierarchies, leading to more stable and discriminative cross-modal representations.
The above methods focus on alleviating the dependency on human-labeled data in the ReID and provide valuable insights for UVGL. However, UVGL suffers from more severe viewpoint and scale variations. Furthermore, the spatial continuity of geographic environments inevitably induces feature ambiguity, severely degrading clustering effectiveness. Thereby, we incorporate instance-level and cluster-level features into a self-supervised UVGL method to learn more discriminative and robust cross-view representations.
Figure 3: The overall pipeline of our method consists of a lightweight backbone, a dual-path contrastive learning strategy, a dynamic hierarchical memory learning module, and an information consistency evolution module. Specifically, the dual-path contrastive learning is designed to learn discriminative and consistent intra-view feature representations. The dynamic hierarchical memory module further captures intra-view feature variations under different perspectives and scales, thereby enhancing the robustness and discriminability of the learned representations. The information consistency evolution module focuses on modeling cross-view feature consistency through a neighborhood-driven learning strategy, and further improves the training process by integrating a pseudo-label enhancement strategy.
@@ III METHOD
As illustrated in Fig. 3, the proposed framework is composed of two key modules, DHML and ICEL. Additionally, a PLE strategy is employed to optimize the self-supervised learning process. Specifically, we adopt a weight-shared ConvNeXt-Tiny network as the backbone, equipped with cluster memory and a dual-path contrastive learning strategy to learn intra-view feature representations. To capture the dynamic variations of intra-class features caused by changes in viewpoint and scale, the DHML module establishes a collaborative relationship between short-term and long-term memory, thereby enhancing the discriminability of intra-view feature representations. Furthermore, the ICEL module leverages instance memory to construct neighborhood-driven dynamic constraints. By integrating neighborhood consistency with mutual information optimization, it guides cross-view features to form unified and consistent representations.
Problem Formulation: In the UVGL, we are given a set of paired UAV-satellite images, denoted by $\{x^{d},x^{s}\}$. Here, $x^{d}$ represents UAV-view images and $x^{s}$ represents satellite-view images. Unlike the supervised setting, our self-supervised method is developed without using paired UAV-satellite images, where the correspondences between UAV-view and satellite-view images are entirely unavailable. Thereby, we randomly sample UAV and satellite images independently during training. The objective is to discover reliable cross-view correspondences under such conditions and accomplish the UVGL task.
III-A Dual-Path Contrastive Learning Baseline
Cluster-based contrastive learning methods have achieved remarkable success in various fields. However, in UVGL, due to the unique challenges of cross-view scenarios, single-path cluster-based contrastive learning struggles to effectively capture view-specific features. To address this issue, we design the baseline network with a dual-path architecture, where UAV-view and satellite-view features are extracted using a shared backbone and clustered to assign pseudo-labels. Subsequently, a dual-path contrastive learning strategy [53] is employed to enhance intra-view representation learning.